OpenStudy (anonymous):

There are 13 patients in Dr. Ziglar's waiting room. Dr. Ziglar can see 7 patients before lunch. In how many different orders can Dr. Ziglar see 7 of the patients before lunch?

OpenStudy (anonymous):

i think its either 13p7 or 13c7 right?

OpenStudy (anonymous):

\[13\times 12\times 11\times 10\times 9\times 8\times 7\]

OpenStudy (anonymous):

seven whole numbers starting at 13 and counting down